OpenStudy (anonymous):

Given the equation x^2-x=0 what constant must be added to both sides so the left side is the square of a binomial?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

what about on x^2+3x=0

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

ok hmm 2.25 i think

OpenStudy (anonymous):

(x + 1.5)^2 = x^2+3x + 2.25 - yep

OpenStudy (anonymous):

alright thanks again mate!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

btw it is easier in these problem to work the fractions than decimals. just say "half of 3 is \[\frac{3}{2}\] and \[(\frac{3}{2})^2=\frac{9}{4}\]"